# Tra-MoE: Learning Trajectory Prediction Model from Multiple Domains for Adaptive Policy Conditioning

### v1

@@ 1 Introduction
In computer vision and natural language processing, significant progress has been made through learning models from multiple domains data. For example, many works use board data to perform self-supervised pre-training, and then adapt to diverse downstream tasks through zero-shot [9, 46, 1] or fine-tuning [13, 11, 20, 55]. Additionally, some research efforts [40, 66, 65, 48] focus on learning a generalist model capable of unifying various tasks. In contrast, in robot learning, due to data scarcity and homogeneity, some groundbreaking methods [34, 2] resort to training only using in-domain data. Recently, some works [49, 3, 58, 4] have shifted to learn from action-free human video data. A particularly scalable and promising paradigm [56, 60, 61, 5] involves learning a trajectory prediction model from action-free video data, and then using small-scale action-labeled demonstrations to learn trajectory-guided policies, thereby achieving greater sample efficiency and better generalization.
@@
  • Extensive simulation and real-world environments demonstrate that our Tra-MoE (with Top-1 gating) can more effectively benefit from large-scale out-of-domain data compared to the dense baseline.
@@ 2 Related Work
Learning from multiple domains in Robot Learning. Scaling laws [27] have fueled substantial advancements in various fields. In robot learning, many recent studies [48, 26, 7, 6] investigate how compute, model size, and training data quantity affect the model performance on robotic manipulation tasks. Based on this, recent works have focused on collecting larger-scale real-world robotic datasets [17, 57, 29], and have successfully trained several robotics foundational models [8, 35, 44, 54, 30, 14, 59] that demonstrate superior transfer capabilities. It is worth noting that these works mainly employ pre-training and fine-tuning paradigm rather than joint co-training paradigm, while focusing on evaluating the combination generalization and neglecting the performance impact on in-domain tasks. Additionally, these works directly combine heterogeneous robotic data with different action and observation spaces for joint training, potentially leading to sub-optimal solutions. In contrast, we leverage sparsely-gated MoE to incorporate larger-scale, out-of-domain, and action-less video data. Furthermore, our experiments demonstrate the importance of MoE for expanding out-of-domain data due to its superior capacity for parameter cooperation and specialization.
@@
Figure 3: (a) The pipeline of our sparsely-gated MoE-based trajectory model (Tra-MoE). (b) The pipeline of our trajectory-guided policy using the adaptive policy conditioning technique.
Representations and Techniques for Policy Conditioning. Policy conditioning equips a multi-task robot policy with the ability to modulate its behavior by incorporating task-specific guidance information. This information, which may take various forms, guides the policy’s action generation, enabling a single model to effectively learn and perform diverse tasks. Classic policy conditioning representations often encompass task identifiers [47], natural language instructions [51, 19, 42, 7], goal images [41], and demonstration videos [24, 22]. These are typically integrated with visual observations within the policy through concatenation or attention mechanisms. Recent works have explored alternative policy conditioning representations to enhance generalization and sample efficiency. These include leveraging internet-trained models to obtain bounding box [52] and object mask [67, 62, 39], as well as learning plan [4], affordance [3, 43], future observations [16, 15, 63, 31], 2D or 3D trajectory [58, 60, 64, 5, 61] from human video data, and also hand-drawn trajectory [18] or sketch [53]. Different from these approaches that directly utilize raw or manually crafted signals for integration with visual observations, we introduce an adaptive policy conditioning technique that explicitly maps 2D trajectories onto the visual observations in space and encodes them as learnable embeddings. These designs ensure adaptive 2D trajectory representation and explicit spatial alignment with the visual observations.
@@ 3.2 Architecture
Based on the aforementioned analysis, we propose an adaptive policy conditioning technique to incorporate trajectory information into 2D images. Specifically, we construct an additional mask modality, and populate it with 2D trajectories based on their specific spatial locations, while setting the value of each point in the trajectory as a learnable embedding. Before encoding the image, we concatenate this mask modality with the image observations on the channel dimension, forming a tensor of dimensions $H\times W\times 4$, as shown in Fig. 3(b). To sum up, our adaptive policy conditioning technique yield an adaptive 2D trajectory representation, which could be explicitly aligned with the image observations to guide the policy prediction more flexibly.
@@ 4 Experiments
We perform evaluation experiments using the LIBERO benchmark [38] and a real low-cost dual-arm robot [32]. In experiments, we qualitatively find that the success rate of downstream manipulation tasks is generally positively correlated with the performance of the trajectory model. Meanwhile, CoTracker [28] can provide relatively accurate trajectory annotations; therefore, we report the average success rate of downstream manipulation tasks as our metric. Our experiments aim to study the following questions:
@@
Q2: Does sparsely-gated Tra-MoE, compared to the dense Tra-baseline, benefit more from large-scale out-of-domain data and demonstrate superior scaling up capabilities?
@@
Table 1: The ablation experiments of the sparsely-gated Tra-MoE. All baselines is the track transformer of ATM.
@@ 4.1 Simulation Experiments
Experiment setup. The LIBERO [38] is divided into five categories: LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, LIBERO-Long, and LIBERO-90. LIBERO-90 includes 90 tasks, while each of the other categories contains 10 tasks. Compared to other categories, LIBERO-90 presents a highly varied selection of object categories, layouts, environments, and skill goals. It is typically used for pre-training rather than joint co-training. Therefore, our main ablation experiments select 90 tasks from LIBERO-90 to construct dataset $\boldsymbol{\mathcal{D}_{ood}}$, utilizing 20 action-free videos per task; and choose 40 tasks from LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long to build dataset $\boldsymbol{\mathcal{D}_{in}}$, using 10 action-labeled demonstrations per task. In summary, for training Tra-MoE, our main ablation experiments use a total of $2,200$ action-free videos, with an out-of-domain to in-domain data ratio of $9:2$. Additionally, on this basis, we also further select 92 tasks from the RLbench [23] simulation environment and generate 5 videos for each task, totaling $2,660$ videos in conjunction with the LIBERO data, to train Tra-MoE. The different physics engines and dynamics properties further increase the data distribution discrepancy in $\boldsymbol{\mathcal{D}_{ood}}$. Finally, as for the trajectory-guided policy, we train a separate multi-task policy for each category. Each policy undergoes behavioral cloning training with 100 action-labeled demonstrations from 10 tasks.
@@
Table 4: The ablation experiments on real-world tasks.
@@
Figure 4: Left: The real-world experiments hardware platform. Right: The real-world evaluation demonstrations.
@@
Result 4: In Tab. 3, we investigate the effectiveness of the adaptive policy conditioning. Specifically, building upon ATM, we first introduce an additional hand-drawn mask modality before image encoding. In this mask, the first half of the trajectory is set to $128$, while the second half is set to $255$. We call it ‘ATM + hand-drawn mask’ and this is similar to RT-Trajectory [18]. Next, we replace the hand-drawn mask with an adaptive mask, where each point on the trajectory is instantiated as a learnable embedding. We call it ‘ATM + adaptive mask’. The results show that ‘ATM + hand-drawn mask’ achieves only a 0.1 improvement compared to the baseline. However, except for a significant performance decrease in LIBERO-Goal, it achieves performance improvements of 6.5, 12.0, and 5.0 in LIBERO-Spatial, LIBERO-Object, and LIBERO-Long respectively. In addition, the ‘ATM + adaptive mask’ achieves a performance improvement of 5.4 (66.8 vs. 61.4) compared to the baseline. It further improves performance across all suites compared to ‘ATM + hand-drawn mask’, especially achieving a 16.0 performance gains in LIBERO-Goal (74.0 vs. 58.0).
@@ 4.2 Real-World Experiments
Experiment setup. As shown in Fig 4, we utilize a dual-arm robot [32] to conduct real-world experiments. We evaluate 5 tasks, which include two single-arm tasks: pouring water and pushing a vegetable to the side of the cutting board, as well as three dual-arm tasks: pulling out tissues, folding towels, and picking up and passing a holder. Specifically, for each task, we collect 50 action-labeled robot demonstrations. Concurrently, we also collect 50 videos of human performing for each task. Compared to the robot data, the human video encompass more environment variations and distractor objects. Please see the appendix for more detailed evaluation details and real-world setup.
@@ A.2 The stability of LIBERO experiment results
A.3 The comparison of trajectory prediction model
In our work, we find that CoTracker [28] can provide highly accurate labels. On the one hand, we can directly use the ground truth provided by CoTracker to calculate the MSE error with our prediction results to measure the performance of trajectory prediction. On the other hand, considering that trajectory prediction task is multi-modal, we also need directly visualize some samples for prediction performance analysis. In our experiments, we qualitatively and quantitatively find that the success rate of downstream manipulation tasks is generally positively correlated with the performance of the trajectory model. Therefore, we report the average success rate of downstream manipulation tasks as our metric.
A.4 The Simulation environments details
A.5 The Real-World environments details

### v2

@@ 1 Introduction
In computer vision and natural language processing, significant progress has been made through learning models from multiple domains data. For example, many works use board data to perform self-supervised pre-training, and then adapt to diverse downstream tasks through zero-shot [10, 49, 1] or fine-tuning [15, 13, 23, 58]. Additionally, some research efforts [43, 72, 71, 51] focus on learning a generalist model capable of unifying various tasks. In contrast, in robot learning, due to data scarcity and homogeneity, some groundbreaking methods [37, 2] resort to training only using in-domain data. Recently, some works [52, 3, 61, 4, 69] have shifted to learn from action-free human video data. A particularly scalable and promising paradigm [59, 63, 65, 5, 20] involves learning a trajectory prediction model from action-free video data, and then using small-scale action-labeled demonstrations to learn trajectory-guided policies, thereby achieving greater sample efficiency and better generalization.
@@
  • Extensive simulation and real-world environments demonstrate that our Tra-MoE (with Top-1 gating) can more effectively benefit from large-scale out-of-domain data (such as cross-embodiments data and different physics engine simulated data) compared to the dense baseline.
@@ 2 Related Work
Learning from multiple domains in Robot Learning. Scaling laws [30] have fueled substantial advancements in various fields. In robot learning, many recent studies [51, 29, 8, 7] investigate how compute, model size, and training data quantity affect the model performance on robotic manipulation tasks. Based on this, recent works have focused on collecting larger-scale real-world robotic datasets [19, 60, 32, 64, 11], and have successfully trained several robotics foundational models [9, 38, 47, 57, 33, 16, 62, 6] that demonstrate superior transfer capabilities. It is worth noting that these works mainly employ pre-training and fine-tuning paradigm rather than joint co-training paradigm, while focusing on evaluating the combination generalization and neglecting the performance impact on in-domain tasks. Additionally, these works directly combine heterogeneous robotic data with different action and observation spaces for joint training, potentially leading to sub-optimal solutions. In contrast, we leverage sparsely-gated MoE to incorporate larger-scale, out-of-domain, and action-less video data. Furthermore, our experiments demonstrate the importance of MoE for expanding out-of-domain data due to its superior capacity for parameter cooperation and specialization.
@@
Figure 3: (a) The pipeline of our sparsely-gated MoE-based trajectory model (Tra-MoE). (b) The pipeline of our trajectory-guided policy using the adaptive policy conditioning technique. Mapping means concatenating the trajectory mask with image observations, while learnable refers to setting each point in the trajectory mask as a learnable embedding.
Representations and Techniques for Policy Conditioning. Policy conditioning equips a multi-task robot policy with the ability to modulate its behavior by incorporating task-specific guidance information. This information, which may take various forms, guides the policy’s action generation, enabling a single model to effectively learn and perform diverse tasks. Classic policy conditioning representations often encompass task identifiers [50], natural language instructions [54, 22, 45, 8], goal images [44], and demonstration videos [27, 25]. These are typically integrated with visual observations within the policy through concatenation or attention mechanisms. Recent works have explored alternative policy conditioning representations to enhance generalization and sample efficiency. These include leveraging internet-trained models to obtain bounding box [55] and object mask [73, 68, 66, 42], as well as learning plan [4], affordance [3, 46], future observations [18, 17, 67, 34], 2D or 3D trajectory [61, 63, 70, 5, 65] from human video data, and also hand-drawn trajectory [21] or sketch [56]. Different from these approaches that directly utilize raw or manually crafted signals for integration with visual observations, we introduce an adaptive policy conditioning technique that explicitly maps 2D trajectories onto the visual observations in space and encodes them as learnable embeddings. These designs ensure adaptive 2D trajectory representation and explicit spatial alignment with the visual observations.
@@ 3.2 Architecture
Based on the aforementioned analysis, we propose an adaptive policy conditioning technique to incorporate trajectory information into 2D images. Specifically, we construct an additional mask modality, and populate it with 2D trajectories based on their specific spatial locations, while setting the value of each point in the trajectory mask as a learnable embedding. Before encoding the image, we concatenate this mask modality with the image observations on the channel dimension, forming a tensor of dimensions $H\times W\times 4$, as shown in Fig. 3(b). To sum up, our adaptive policy conditioning technique yield an adaptive 2D trajectory representation, which could be explicitly aligned with the image observations to guide the policy prediction more flexibly.
@@ 4 Experiments
We perform evaluation experiments using the LIBERO benchmark [41] and a real low-cost dual-arm robot [35]. In experiments, we quantitatively find that the success rate of downstream manipulation tasks is generally positively correlated with the performance of the trajectory model. Detailed quantitative experiment results can be found in the appendix. Meanwhile, CoTracker [31] can provide relatively accurate trajectory annotations; therefore, we report the average success rate of downstream manipulation tasks as our metric. Our experiments aim to study the following questions:
@@
Q2: Does sparsely-gated Tra-MoE, compared to the dense Tra-baseline, benefit more from large-scale out-of-domain data (such as cross-embodiments and different physics engine data) and demonstrate superior scaling up capabilities?
@@
Table 1: The ablation experiments of the sparsely-gated Tra-MoE. All baselines are the track transformer of ATM.
@@ 4.1 Simulation Experiments
Experiment setup. The LIBERO [41] is divided into five categories: LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, LIBERO-Long, and LIBERO-90. LIBERO-90 includes 90 tasks, while each of the other categories contains 10 tasks. Compared to other categories, LIBERO-90 presents a highly varied selection of object categories, layouts, environments, and skill goals. It is typically used for pre-training rather than joint co-training. Therefore, our main ablation experiments select 90 tasks from LIBERO-90 to construct dataset $\boldsymbol{\mathcal{D}_{ood}}$, utilizing 20 action-free videos per task; and choose 40 tasks from LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long to build dataset $\boldsymbol{\mathcal{D}_{in}}$, using 10 action-labeled demonstrations per task. In summary, for training Tra-MoE, our main ablation experiments use a total of $2,200$ action-free videos, with an out-of-domain to in-domain data ratio of $9:2$. Additionally, on this basis, we also further select 92 tasks from the RLbench [26] (using CoppeliaSim rather than MuJoCo as physics engine) simulation environment and generate 5 videos for each task, totaling $2,660$ videos in conjunction with the LIBERO data, to train Tra-MoE. The different physics engines and dynamics properties further increase the data distribution discrepancy in $\boldsymbol{\mathcal{D}_{ood}}$. Finally, as for the trajectory-guided policy, we train a separate multi-task policy for each category. Each policy undergoes behavioral cloning training with 100 action-labeled demonstrations from 10 tasks.
@@
Table 4: The ablation of ood human data, MoE architecture, and adaptive policy conditioning technique on our real-world tasks.
@@
Figure 4: Left: The real-world experiments hardware platform setup. Right: The real-world tasks evaluation demonstrations.
@@
Result 4: In Tab. 3, we investigate the effectiveness of the adaptive policy conditioning. Specifically, building upon ATM, we first introduce an additional hand-drawn mask modality before image encoding. In this mask, the first half of the trajectory is set to $128$, while the second half is set to $255$. We call it ‘ATM + hand-drawn mask’ and this is similar to RT-Trajectory [21]. Next, we replace the hand-drawn mask with an adaptive mask, where each point on the trajectory is instantiated as a learnable embedding. We call it ‘ATM + adaptive mask’. The results show that ‘ATM + hand-drawn mask’ achieves only a 0.1 improvement compared to the baseline. However, except for a significant performance decrease in LIBERO-Goal, it achieves performance improvements of 6.5, 12.0, and 5.0 in LIBERO-Spatial, LIBERO-Object, and LIBERO-Long respectively. In addition, the ‘ATM + adaptive mask’ achieves a performance improvement of 4.9 (66.3 vs. 61.4) compared to the baseline. It further improves performance across all suites compared to ‘ATM + hand-drawn mask’, especially achieving a 19.0 performance gains in LIBERO-Goal (77.0 vs. 58.0).
@@ 4.2 Real-World Experiments
Experiment setup. As shown in Fig 4, we utilize a dual-arm robot [35] to conduct real-world experiments. We evaluate 5 tasks, which include two single-arm tasks: pouring water and pushing a vegetable to the side of the cutting board, as well as three dual-arm tasks: pulling out tissues, folding towels, and picking up and passing a holder. Specifically, for each task, we collect 50 action-labeled robot demonstrations. Concurrently, we also collect 50 videos of human (cross-embodiments) performing for each task. Compared to the robot data, the human video encompass more environment variations and distractor objects. Please see the appendix for more detailed evaluation details and real-world setup.
@@ 5 Conclusion
Acknowledgements
This work is supported by the National Key R&\&&D Program of China (No. 2022ZD0160900), Jiangsu Frontier Technology Research and Development Program (No. BF2024076), and the Collaborative Innovation Center of Novel Software Technology and Industrialization.
Supplementary Material
@@ A.2 The stability of LIBERO experiment results
A.3 The comparison of trajectory prediction model
In our work, we find that CoTracker [31] can provide highly accurate labels. On the one hand, we can directly use the ground truth provided by CoTracker to calculate the MSE error with our prediction results to measure the performance of trajectory prediction. On the other hand, considering that trajectory prediction task is multi-modal, we also need directly visualize some samples for prediction performance analysis. In our experiments, we measure trajectory prediction performance using the MSE loss on validation set of four evaluation suites, as well as qualitatively and quantitatively find that the success rate of downstream manipulation tasks is generally positively correlated with the performance of the trajectory model, as shown in Fig. 5. Therefore, we report the average success rate of downstream manipulation tasks as our metric.
Figure 5: The quantitative relationship between downstream policy success rate and trajectory model performance.
@@ A.3 The comparison of trajectory prediction model
A.4 The Simulation environments details
A.5 The Real-World environments details